Find the products AB and BA for the diagonal matrices. -=[ -), 0-105] AB = BA =
Answer #1

A = [ 2 0 7 10-3 J B = -4 07 Io5] AB= [20] 1-4 07 = [20-4)+00) LO 3 |Los ) 10(-4) + (-3)0) 2(0)+015) 0(0)+63X5) - AB =[-8 OL